OpenStudy (anonymous):

Hi, how may solve this equation? 2-x=10

OpenStudy (anonymous):

Bring constants to one side and solve for x.

OpenStudy (anonymous):

by solving for x. and as GT is directing u

OpenStudy (saifoo.khan):

2 -x = 10 -x = 10 - 2 -x = 8 x = -8

OpenStudy (anonymous):

so: 2-x=10 2-2-x=10-2 -x=8 -1*-x=-1*8 x=-8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

for this answer saifoo one of ur medal is snatched away from u

OpenStudy (saifoo.khan):

LOL. that's cruel. :'(

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

yes it is but don't do others work these guys are taking us as a medium of getting their work done

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

take a serious note of this fact no1 here is ready to understand a thing they just want readymade answers

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

That is the most disappointing thing about this site. When I first looked at it, I was excited about the concept.

OpenStudy (anonymous):

Now I know it is mostly used for getting answers.

OpenStudy (anonymous):

I appreciate your help. I forgot algebra. I was trying to help my daughter so that I can help her.

OpenStudy (anonymous):

gamze - when someone gives an answer, insist on giving an explanation as to how the answer was obtained and how similar problems can be solved with no more help in the future. Otherwise, it is a waste of time.

OpenStudy (anonymous):

I think I bothered some people in this group

OpenStudy (anonymous):

I will keep asking questions as far as someone answers.

OpenStudy (anonymous):

x=-8 :)